# On the transitivity of Anosov diffeomorphisms

### v2

@@ 3. Examples
We start with an example of a homeomorphism on a compact metric space which has the $su-$intersecting property in the whole domain but fails to satisfy both the two-sided limit shadowing and the shadowing property. Consider a two-sided shift $\sigma$ on a space with a finite alphabet endowed with the usual metric which has the specification property but is not a shift of finite type. Then $\sigma$ has the $su-$intersecting property in the whole space (cf. [15][Proposition 5]), even though $\sigma$ fails to satisfy both the two-sided limit shadowing property (cf. [15][Corollary 6] and the shadowing property (cf. [62]).
@@
According to [24, Theorem 1.3, Corollary 1.4], a generic element in $\operatorname{Hom}(X,\mu_{0})$ is topologically mixing and satisfies the shadowing property. Using Theorem A(a), we conclude that, $C^{0}$ generically in $\operatorname{Hom}(X,\mu_{0})$, a homeomorphism has the barycenter property in $A$ for every nonempty subset $A$ of $X$.
@@
The marix $A$ is irreducible (to check this it is enough to compute $A^{2}$ and $A^{3}$) but it is not aperiodic. Therefore, the homeomorphism $\sigma_{|\Sigma_{A}}\colon\Sigma_{A}\to\Sigma_{A}$ is a shift of finite type (hence it has the shadowing property), is topologically transitive (cf. [63, Theorem 1.19]), but is not topologically mixing (cf. [63, Theorem 1.31]). Moreover, the space $\Sigma_{A}$ is perfect (that is, it is closed, totally disconnected and has no isolated points) due to the fact that the matrix $A$ has lines whose entries sum is equal to $2$ (cf. [50, Proposition 2.8]). Therefore, Theorem A(a) indicates that $\sigma_{|\Sigma_{A}}$ has the barycenter property in $\Sigma_{A}$, whereas Theorem A(c) shows that $\sigma_{|\Sigma_{A}}$ fails to satisfy the $su-$intersecting property in $\Sigma_{A}$.
@@
The next example is mentioned in [34, Remark 1.1] to justify that the barycenter property in $\operatorname{Per}(f)$, though a kind of tracing condition, is ultimately distinct from the specification property. Let $f\colon S^{1}\to S^{1}$ be a $C^{1}$ diffeomorphism whose non-wandering set reduces to two fixed points which are saddle-nodes. Then $f$ does not have the shadowing property (cf. [48, Theorem 3.1.1]) and is not expansive (cf. [63, Theorem 5.27]). Moreover, $f$ is not topologically transitive, so it does not have the specification property (cf. [18, Proposition 21.3]. However, $f$ has the $su-$intersecting property in $\operatorname{Per}(f)$, hence the barycenter property in $\operatorname{Per}(f)$.
We observe that, on the contrary, if $T\colon X\to X$ is a homeomorphism of a compact metric space $(X,d)$ without isolated points, $A$ is a nonempty subset of $X$ and $T$ has the specification property, then $T$ has the barycenter property in $A$. Let us see why. Take $p,\,q\in A$ and fix $\varepsilon>0$. Since $T$ has the specification property there exists $N=N(\varepsilon)\in\mathbb{N}$ such that any $N-$spaced specification $\xi$ is $\varepsilon-$traced by some $y\in X$. Given $n_{1},n_{2}\in\mathbb{N}$, define
@@
A Morse-Smale diffeomorphism of the unit circle is a simple example of an $\Omega-$stable diffeomorphism of a smooth closed Riemannian manifold which has the shadowing property (cf. [48]), is not expansive (cf. [63, Theorem 5.27]) and does not satisfy the barycenter property in the set of periodic points (as easily concluded from Theorem A(b)).
@@
Claim 1: The identity map on a compact metric space $(X,d)$ has the shadowing property if and only if $X$ is totally disconnected (cf. [6][Theorem 2.3.2]).
@@
Claim 3: The presence of the $su-$intersecting property induces some rigidity on the dynamics.
@@
Let $T\colon X\to X$ be a homeomorphism of a compact metric space $(X,d)$ without isolated points. If $T$ has the $su-$intersecting property in $X$ then $T$ is chain-transitive.
Proof.
See [14][Lemma 3.6] or [33][Lemma 4.2]. ∎
@@
    If $T^{k}$ is chain-transitive for every positive integer $k$, then $T$ does not have the shadowing property.
@@
Consequently, the identity map on a totally disconnected compact metric space $X$, containing at least two elements and without isolated points, has the shadowing property but does not satisfy the $su-$intersecting property in $X$, nor the barycenter property in $X$ according to Theorem A(b).
@@
Therefore, for any $n_{1},n_{2}\in\mathbb{N}$, one can find an integer $m=N\in[0,N]$ and $x_{0}\in X$ satisfying $d\big{(}T^{i}(x_{0}),T^{i}(p)\big{)}<\varepsilon$ for every $-n_{1}\leqslant i\leqslant 0$ and $d\big{(}T^{i+m}(x_{0}),T^{i}(q)\big{)}<\varepsilon$ for every $0\leqslant i\leqslant n_{2}$. Thus, the pair $p,q$ complies with the condition of the barycenter property in $\operatorname{Per}(T)$. However, if we exchange the fixed points this condition is no longer valid. Thus, $T$ does not have the barycenter property in $\operatorname{Per}(T)$. We note that $\Omega(T)=\overline{\operatorname{Per}(T)}$, $T$ has the shadowing property (see [46, Proposition 3]) and is topologically transitive. However, $X$ has isolated points and the conclusion of Theorem A(a) does not hold.
@@
Both Theorems A and B assume that the dynamics has the shadowing property. The next example indicates that this is an essential hypothesis to obtain topological transitivity.
@@ 4. Proof of Theorem A
In this section we explain why, under the assumptions of Theorem A, the barycenter property in a nonempty subset $A$ such that $T(A)=A$ and $\Omega(T)\subseteq\overline{A}$ yields topological transitivity.
$(a)$ Let $T\colon X\to X$ be a homeomorphism of a compact metric space $(X,d)$ without isolated points and consider a nonempty set $A\subseteq X$. Suppose that $T$ has the shadowing property and is topologically transitive. We will prove that $T$ has the barycenter property in $A$.
Take $p,\,q\in A$ and fix $\varepsilon>0$. Let $\delta>0$, provided by the shadowing property, be such that every $\delta-$pseudo-orbit of $T$ is $\varepsilon-$shadowed in $X$. Consider $U=B(T(p),\delta)$ and $V=B(q,\delta)$, where $B(a,r)$ stands for the open ball in $(X,d)$ centered at $a$ with radius $r$. From [4][Theorem 1.4], one has that one-sided topological transitivity is equivalent to topological transitivity whenever the compact metric space does not contain isolated points. So, $T$ is one-sided topologically transitive. Hence, there exists $m\in\mathbb{N}$ such that $T^{m}(U)\cap V\neq\emptyset$. Select $x\in B(T(p),\delta)$ such that $T^{m}(x)\in B(q,\delta)$, and define a $\delta-$pseudo-orbit as follows:
$x_{i}$ $=$ $T^{i}(p)\quad\quad\forall\,i\leqslant 0$
$x_{1}$ $=$ $x,\quad x_{2}\,=\,T(x),\quad\cdots,\quad x_{m}\,=\,T^{m-1}(x)$
$x_{(m+1)+\ell}$ $=$ $T^{\ell}(q)\quad\quad\forall\,\ell\geqslant 0.$
By the shadowing property, there exists $y\in X$ such that $d(T^{k}(y),x_{k})<\varepsilon$ for all $k\in\mathbb{Z}$. In particular,
$d(T^{i}(y),\,x_{i})$ $=$ $d(T^{i}(y),\,T^{i}(p))\,<\,\varepsilon\quad\quad\forall\,i\leqslant 0$
$d(T^{(m+1)+\ell}(y),\,x_{(m+1)+\ell})$ $=$ $d(T^{(m+1)+\ell}(y),\,T^{\ell}(q))\,<\,\varepsilon\quad\quad\forall\,\ell\geqslant 0.$
Therefore, $T$ has the barycenter property in $A$. The proof of item $(a)$ of Theorem A is complete.
$(b)$ Consider a homeomorphism $T\colon X\to X$ of a compact metric space $(X,d)$ without isolated points such that $T$ has the shadowing property, $T(A)=A$ and $\Omega(T)\subseteq\overline{A}$. Assume that $T$ satisfies the barycenter property in $A$.
@@
Let $T\colon X\to X$ be a homeomorphism of a compact metric space $(X,d)$. Assume that $T(A)=A$, $\Omega(T)\subseteq\overline{A}$ and that $T$ has the barycenter property in $A$. Then $T$ is chain-transitive.
@@
Since $\Omega(T)\subseteq\overline{A}$ and $a,b\in\Omega(T)$, there are $p,\,q\in A$ such that
@@
Summoning $\varepsilon>0$ and the barycenter property for $T(p)$ and $T^{-1}(q)$, which is possible because $T(A)=A$, we take a positive integer $N=N(\varepsilon,T(p),T^{-1}(q))$ such that, for any $n_{1},n_{2}\in\mathbb{N}$, one can find an integer $m\in[0,N]$ and $z\in X$ satisfying
@@
Let us resume the proof of Theorem A(b). By Lemma 4.1, $T$ is chain-transitive. As, by assumption, $T$ also has the shadowing property, then $T$ is one-sided topologically transitive by [31][Lemma 2.1]. Thus, $T$ is topologically transitive. The proof of Theorem A(b) is complete.
$(c)$ Let $T\colon X\to X$ be a homeomorphism, of a compact metric space $(X,d)$ without isolated points, which satisfies the shadowing property and is accessible in $R(T)$. We start by showing the following auxiliar result which replaces, in this context, Lemma 4.1.
@@
We may now finish the proof of Theorem A(c). From [17][Corollary 2.3], since $T$ has the shadowing property we know that $\overline{R(T)}=\Omega(T)$. On the other hand, by Lemma 4.2 $T$ is chain-transitive. As, by assumption, $T$ also has the shadowing property, then $T$ is one-sided topologically transitive due to [31][Lemma 2.1]. Thus, $T$ is topologically transitive.
$(d)$ Let $T\colon X\to X$ be a homeomorphism, of a compact metric space $(X,d)$ without isolated points, which satisfies the shadowing property and has the $su-$intersecting property in $X$.
By Lemma 3.1, we know that, for every positive integer $k$, the homeomorphism $T^{k}$ has the $su-$intersecting property in $X$. According to Lemma 3.2, we also have that $T^{k}$ is chain-transitive for every positive integer $k$. Moreover, as $T$ has the shadowing property then $T^{k}$ has the shadowing property for every positive integer $k$ (cf. [6][Theorem 2.3.3]). Since the shadowing property together with the chain-transitivity imply topological transitivity, we conclude that $T^{k}$ is topologically transitive for every positive integer $k$. Finally, by [30][Theorem 45], since $T$ has the shadowing property and $T^{k}$ is topologically transitive for every positive integer $k$, then $T$ is topologically mixing. This ends the proof of Theorem A.

### v3

@@ 3. Examples
We start with an example of a homeomorphism on a compact metric space which has the $su-$intersecting property in the whole domain but fails to satisfy both the two-sided limit shadowing and the shadowing property. Consider a two-sided shift $\sigma$ on a space with a finite alphabet endowed with the usual metric which has the specification property but is not a shift of finite type. Then $\sigma$ has the $su-$intersecting property in the whole space (cf. [15, Proposition 5]), even though $\sigma$ fails to satisfy both the two-sided limit shadowing property (cf. [15, Corollary 6] and the shadowing property (cf. [61]).
@@
According to [24, Theorem 1.3, Corollary 1.4], a generic element in $\operatorname{Hom}(X,\mu_{0})$ is topologically mixing and satisfies the shadowing property. Using Theorem A(c), we conclude that, $C^{0}$ generically in $\operatorname{Hom}(X,\mu_{0})$, a homeomorphism has the barycenter property in $A$ for every nonempty subset $A$ of $X$.
@@
The matrix $A$ is irreducible (to check this it is enough to compute $A^{2}$ and $A^{3}$) but it is not aperiodic. Therefore, the homeomorphism $\sigma_{|\Sigma_{A}}\colon\Sigma_{A}\to\Sigma_{A}$ is a shift of finite type (hence it has the shadowing property), is topologically transitive (cf. [62, Theorem 1.19]), but is not topologically mixing (cf. [62, Theorem 1.31]). Moreover, the space $\Sigma_{A}$ is perfect (that is, it is closed, totally disconnected and has no isolated points) due to the fact that the matrix $A$ has lines whose entries sum is equal to $2$ (cf. [50, Proposition 2.8]). Therefore, Theorem A(c) indicates that $\sigma_{|\Sigma_{A}}$ has the barycenter property in $\Sigma_{A}$, whereas Theorem A(e) indicates that $\sigma_{|\Sigma_{A}}$ fails to satisfy the $su-$intersecting property in $\Sigma_{A}$.
@@
The next example is mentioned in [34, Remark 1.1] to justify that the barycenter property in $\operatorname{Per}(f)$, though a kind of tracing condition, is ultimately distinct from the specification property. Let $f\colon S^{1}\to S^{1}$ be a $C^{1}$ diffeomorphism whose non-wandering set reduces to two fixed points which are saddle-nodes. Then $f$ does not have the shadowing property (cf. [48, Theorem 3.1.1]) and is not expansive (cf. [62, Theorem 5.27]). Moreover, $f$ is not topologically transitive, so it does not have the specification property (cf. [18, Proposition 21.3]. However, $f$ has the $su-$intersecting property in $\operatorname{Per}(f)$, hence the barycenter property in $\operatorname{Per}(f)$.
We observe that, on the contrary, if $T\colon X\to X$ is a homeomorphism of a compact metric space $(X,d)$ without isolated points, $A$ is a nonempty subset of $X$ and $T$ has the specification property, then $T$ has the barycenter property in $A$. Indeed, take $p,\,q\in A$ and fix $\varepsilon>0$. Since $T$ has the specification property there exists $N=N(\varepsilon)\in\mathbb{N}$ such that any $N-$spaced specification $\xi$ is $\varepsilon-$traced by some $y\in X$. Given $n_{1},n_{2}\in\mathbb{N}$, define
@@
A Morse-Smale diffeomorphism of the unit circle is a simple example of an $\Omega-$stable diffeomorphism of a smooth closed Riemannian manifold which has the shadowing property (cf. [48]), is not expansive (cf. [62, Theorem 5.27]) and does not satisfy the barycenter property in the set of periodic points (as is easily concluded from Theorem A(d)).
@@
Claim 1: The identity map on a compact metric space $(X,d)$ has the shadowing property if and only if $X$ is totally disconnected (cf. [6, Theorem 2.3.2]).
@@
Claim 3: The $su-$intersecting property induces some rigidity on the dynamics.
@@
[14, Lemma 3.6] or [33, Lemma 4.2] Let $T\colon X\to X$ be a homeomorphism of a compact metric space $(X,d)$ without isolated points. If $T$ has the $su-$intersecting property in $X$ then $T$ is chain-transitive.
@@
    If $T^{k}$ is chain-transitive for every $k\in\mathbb{N}$, then $T$ does not have the shadowing property.
@@
Consequently, the identity map on a totally disconnected compact metric space $X$, containing at least two elements and without isolated points, has the shadowing property but does not satisfy the $su-$intersecting property in $X$, nor the barycenter property in $X$ according to Theorem A(d).
@@
Therefore, for any $n_{1},n_{2}\in\mathbb{N}$, one can find an integer $m=N\in[0,N]$ and $x_{0}\in X$ satisfying $d\big{(}T^{i}(x_{0}),T^{i}(p)\big{)}<\varepsilon$ for every $-n_{1}\leqslant i\leqslant 0$ and $d\big{(}T^{i+m}(x_{0}),T^{i}(q)\big{)}<\varepsilon$ for every $0\leqslant i\leqslant n_{2}$. Thus, the pair $p,q$ complies with the condition of the barycenter property in $\operatorname{Per}(T)$. However, if we exchange the fixed points this condition is no longer valid. Thus, $T$ does not have the barycenter property in $\operatorname{Per}(T)$. We note that $\Omega(T)=\overline{\operatorname{Per}(T)}$, $T$ has the shadowing property (see [46, Proposition 3]) and is topologically transitive. However, $X$ has isolated points and the conclusion of Theorem A(c) does not hold.
@@
Both Theorems A and B assume that the homeomorphism has the shadowing property. The next example indicates that this is an essential hypothesis to obtain topological transitivity.
@@ 4. Proof of Theorem A
In this section we shall explain why, under the assumptions of Theorem A, the topological transitivity implies that the barycenter property is valid in any nonempty subset $A$ of $X$; and, conversely, for what reason the barycenter property in an invariant subset $A$ such that $\Omega(T)=\overline{A}$ yields topological transitivity.
Let $T\colon X\to X$ be a homeomorphism of a compact metric space $(X,d)$ without isolated points.
(a) Suppose that $A$ is dense in $X$ and $T$ has the barycenter property in $A$. Let $U$ and $V$ be nonempty open subsets of $X$. Since $A$ is dense in $X$, there are $p\in U$ and $q\in V$ such that $p,q\in A$; moreover, there exists $\varepsilon>0$ such that $B(p,\varepsilon)\subseteq U$ and $B(q,\varepsilon)\subseteq V$. By the barycenter property, there is a positive integer $N=N(\varepsilon,p,q)$ such that, for $n_{1}=1=n_{2}$, one can find an integer $m\in[0,N]$ and $x_{0}\in X$ satisfying $d\big{(}T^{i}(x_{0}),T^{i}(p)\big{)}<\varepsilon$ for every $-n_{1}\leqslant i\leqslant 0$ and $d\big{(}T^{i+m}(x_{0}),T^{i}(q)\big{)}<\varepsilon$ for every $0\leqslant i\leqslant n_{2}$. In particular, $d\big{(}x_{0},p\big{)}<\varepsilon$ and $d\big{(}T^{m}(x_{0}),q\big{)}<\varepsilon$. So, $T^{m}(U)\cap V\neq\emptyset$. Hence $T$ is topologically transitive.
(b) Suppose that $\Omega(T)=CR(T)$ and that there is a subset $A$ of $X$ such that $T(A)=A$, $\Omega(T)=\overline{A}$ and $T$ has the barycenter property in $A$. We start by showing that, under these assumptions, $T$ is chain-transitive.
@@
Let $T\colon X\to X$ be a homeomorphism of a compact metric space $(X,d)$. If there is a subset $A$ of $X$ such that $T(A)=A$, $\Omega(T)=\overline{A}$ and $T$ has the barycenter property in $A$, then $T$ is chain-transitive.
@@
Since $\Omega(T)=\overline{A}$ and $a,b\in\Omega(T)$, there are $p,\,q\in A$ such that
@@
Summoning the barycenter property for $T(p)$ and $T^{-1}(q)$, which is possible because $T(A)=A$, we may take a positive integer $N=N(\varepsilon,T(p),T^{-1}(q))$ such that, for any $n_{1},n_{2}\in\mathbb{N}$, one can find an integer $m\in[0,N]$ and $z\in X$ satisfying
@@
By Lemma 4.1, one has that $T$ is chain-transitive, so $CR(T)=X$. Thus, $\Omega(T)=X$ since, by assumption, $\Omega(T)=CR(T)$. From $\Omega(T)=\overline{A}$ and $\Omega(T)=X$, we obtain that $A$ is dense in $X$. So the hypothesis on item (a) of Theorem A are fulfilled, and therefore $T$ is topologically transitive.
(c) Suppose that $T$ is topologically transitive and has the shadowing property. The following auxiliary result provides the maximum length of a chain between any two elements of $X$ when they are subject to a topologically transitive dynamics.
Lemma 4.2.
Let $T\colon X\to X$ be a homeomorphism of a compact metric space $(X,d)$ without isolated points. If $T$ is topologically transitive, then for every $\xi>0$ there exists a positive integer $N=N(\xi)$ such that for any $a,b\in X$ there is a $\xi-$chain from $a$ to $b$ with length $L$, where $2\leq L\leq N$.
Proof.
Recall that one-sided topological transitivity is equivalent to topological transitivity whenever the compact metric space does not contain isolated points (cf. [4, Theorem 1.4]). So, $T$ is one-sided topologically transitive. Consider an open finite subcover of $X$, given by $\{U_{1}\cdots,U_{s}\}$, such that $\operatorname{diam}(U_{i})<\xi$ for every $i\in\{1,\cdots,s\}$. Thus, by the one-sided transitivity, for each pair $(i,j)\in\{1,\cdots,s\}\times\{1,\cdots,s\}$ there exists $t_{(i,j)}\in\mathbb{N}$ such that $T^{-t_{(i,j)}}(U_{i})\cap U_{j}\neq\emptyset$. Define
[…]
We claim that for any $a,b\in X$ there is a $\xi-$chain from $a$ to $b$ of length $p$ with $2\leq p\leq N$. Indeed, since $X=\bigcup_{i=1}^{s}U_{i}$, there are $r,m$ such that $T(a)\in U_{r}$ and $b\in U_{m}$. Using that $T^{-t_{(m,r)}}(U_{m})\cap U_{r}\neq\emptyset$, take $c\in U_{r}$ such that $T^{t_{(m,r)}}(c)\in U_{m}$. Then, define a $\xi-$pseudo-orbit as follows:
[…]
Note that the length of this chain is $L=t_{(m,r)}+2$; hence $2\leq L\leq N$. ∎
Let $A$ be a nonempty subset of $X$. Take $p,\,q\in A$ and fix $\varepsilon>0$. Let $\delta>0$, provided by the shadowing property, be such that every $\delta-$pseudo-orbit of $T$ is $\varepsilon-$shadowed in $X$. Consider $U=B(T(p),\delta)$ and $V=B(q,\delta)$, where $B(a,r)$ stands for the open ball in $(X,d)$ centered at $a$ with radius $r$. From [4, Theorem 1.4], we know that one-sided topological transitivity is equivalent to topological transitivity whenever the compact metric space does not contain isolated points. So, $T$ is one-sided topologically transitive. Applying Lemma 4.2 when $\xi=\delta$, we find a positive integer $N=N(\delta)$ such that there is a $\delta-$chain $(y_{j})_{1}^{m}$ from $T(p)$ to $q$ with length $m$, where $2\leq m\leq N$, satisfying
$y_{1}\,=\,T(p)\quad\quad\text{and}\quad\quad y_{m}\,=\ q.$
Define a $\delta-$pseudo-orbit as follows:
$x_{i}$ $=$ $T^{i}(p)\quad\quad\forall\,i\leqslant 0$
$x_{1}$ $=$ $y_{1},\quad x_{2}\,=\,y_{2},\quad\cdots,\quad x_{m-1}\,=\,y_{m-1}$
$x_{m+\ell}$ $=$ $T^{\ell}(q)\quad\quad\forall\,\ell\geqslant 0.$
By the shadowing property, there exists $z\in X$ such that $d(T^{k}(z),x_{k})<\varepsilon$ for all $k\in\mathbb{Z}$. In particular,
$d(T^{i}(z),\,x_{i})$ $=$ $d(T^{i}(z),\,T^{i}(p))\,<\,\varepsilon\quad\quad\forall\,i\leqslant 0$
$d(T^{m+\ell}(z),\,x_{m+\ell})$ $=$ $d(T^{m+\ell}(z),\,T^{\ell}(q))\,<\,\varepsilon\quad\quad\forall\,\ell\geqslant 0.$
Therefore, $T$ has the barycenter property in $A$. ∎
(d) Suppose that $T$ has the shadowing property. Thus, $\Omega(T)=CR(T)$ (cf. [6, Theorem 3.1.2]). Therefore, if there exists $A\subseteq X$ such that $T(A)=A$, $\Omega(T)=\overline{A}$ and $T$ has the barycenter property in $A$, then by item (b) we know that $T$ is topologically transitive. ∎
(e) Assume that $T$ has the shadowing property and is accessible in $R(T)$. We start by proving the following auxiliar result which replaces, in this context, Lemma 4.1.
@@
We may now finish the proof of Theorem A(e). From [17, Corollary 2.3], since $T$ has the shadowing property we know that $\overline{R(T)}=\Omega(T)$. On the other hand, by Lemma 4.3, $T$ is chain-transitive. As, by assumption, $T$ also has the shadowing property, then $T$ is one-sided topologically transitive due to [31, Lemma 2.1]. Thus, $T$ is topologically transitive. ∎
(f) Assume that $T$ has the shadowing property and the $su-$intersecting property in $X$. By Lemma 3.1, we know that, for every positive integer $k$, the homeomorphism $T^{k}$ has the $su-$intersecting property in $X$. Moreover, according to Lemma 3.2, $T^{k}$ is chain-transitive for every positive integer $k$. Besides, as $T$ has the shadowing property then $T^{k}$ also has the shadowing property for every positive integer $k$ (cf. [6, Theorem 2.3.3]). Since the shadowing property together with the chain-transitivity imply topological transitivity, we conclude that $T^{k}$ is topologically transitive for every positive integer $k$. Finally, since $T$ has the shadowing property and $T^{k}$ is topologically transitive for every positive integer $k$, then $T$ is topologically mixing by [30, Theorem 45]. ∎